nightshade0773:

Gwen is running the Challenge of Champions obstacle race at her summer camp. She runs 1/2 of a mile, but then she twists her ankle while going up the Mounds of Doom. Gwen still has 3/4 of a mile left to run, so she drops out of the race

mhanifa:

What is the question? Total distance?

nightshade0773:

yea

mhanifa:

Any idea how to get it?

nightshade0773:

-_- NO

mhanifa:

Sum the two

mhanifa:

Can you?

mhanifa:

\[\frac{ 1 }{2 } + \frac{ 3 }{4 } =\] Can you continue?

mhanifa:

Find the common denominator and add up

mhanifa:

\[\frac{ 1 }{ 2 } = \frac{ 2 }{4 } \] Now, can you continue

mhanifa:

Ok, \[\frac{ 2 }{4 }+\frac{ 3 }{4 } = \frac{ 5}{4 } = 1 \frac{ 1 }{4 } \] miles